# Self-NPO: Negative Preference Optimization of Diffusion Models by Simply Learning from Itself without Explicit Preference Annotations

### v1

@@ 4.2 Comparison
Quantitative comparison. For text-to-image generation, we evaluate our method quantitatively by adhering to prior research and utilizing the ‘test_unique’ split from Pick-a-pic as the benchmark for testing (kirstain2023pick, ). We utilize PickScore (kirstain2023pick, ), HPSv2 (wu2023human, ), ImageReward (xu2024imagereward, ), and Laion-Aesthetic (schuhmann2022laion, ) as our evaluation metrics. The outcomes of this quantitative assessment are detailed in Tables 1 and 7. These tables reveal that integrating Self-NPO with the base model and its preference-optimized variants consistently improves the aesthetic quality of the generated images. Beyond presenting average scores, as shown in Fig. 8, we also compute the percentage of samples generated from the same prompt that attain a higher preference score. The results produced with Self-NPO markedly surpass those generated without it. For text-to-video generation, we follow the experimental setup of VADER prabhudesai2024video but use the more advanced text-to-video diffusion model, CogVideoX-5B, as the baseline. We evaluate performance improvements by employing various prompt sets that focus on different topics, specifically selecting the prompt sets provided by VADER. We report the quantitative evaluation and winning ratios in Table 5. In most cases, Self-NPO consistently enhances both visual quality and aesthetic scores.
Qualitative comparison. We present the generated results in different scenarios in Fig. 1, Fig. 3, Fig. 5, and Fig. 9, demonstrating that self-NPO consistently improves both the generation visual quality and fidelity.
User study. We evaluate generation quality across four areas: Color and Lighting, High-Frequency Details, Low-Frequency Composition, and Text-Image Alignment. In Color and Lighting, users assess whether the images feature natural, visually pleasing color schemes and lighting. For High-Frequency Details, users look at the sharpness and texture details, especially edges and fine elements. In Low-Frequency Composition, they focus on the overall structure and balance of the image. Finally, for Text-Image Alignment, users judge how well the generated image matches the input text. The user study uses prompts from Pickapic ‘validation_unique‘, with different models generating images based on the same random seed. Users choose between "No Preference," (Draws) "Self-NPO is better," (Wins) or "Self-NPO is worse" (Losses) for each pair of images. We collect responses from 9 volunteers, each evaluating 100 pairs of images generated by Dreamshaper, totaling 900 votes. Results, shown in Fig. 7, reveal that NPO significantly improves high-frequency details, enhances color and lighting preferences, and helps with composition. It also shows an improvement in text-image alignment.
4.3 Applications.
@@
Controllable generation. Moreover, our approach seamlessly integrates with controllable generation techniques. By incorporating controllable plugins like T2I-Adapter mou2024t2i , we significantly boost the model’s generation quality under layout-controlled conditions.
@@
Table 1: Quantitative performance comparison with stable diffusion v1-5 based models. means the metrics are copied from NPO papers. Other metrics are tested with official weights.
Method PickScore HPSv2.1 ImageReward Aesthetic
@@
SePPO + Self NPO 21.73 30.28 0.6744 6.014
Diff.-DPO + NPO (reg$=500$) 21.58 27.60 0.3101 5.762
Diff.-DPO + NPO (reg$=1000$) 21.43 27.36 0.3472 5.773
Diff.-DPO + Self-NPO 21.34 27.78 0.4085 5.710

### v2

@@ Introduction
Quantitative comparison. For text-to-image generation, we evaluate our method quantitatively by adhering to prior research and utilizing the ‘test_unique’ split from Pick-a-pic as the benchmark for testing (kirstain2023pick). We utilize PickScore (kirstain2023pick) (P.S.), HPSv2.1 (wu2023human) (HPS.), ImageReward (xu2024imagereward) (I.R.), and Laion-Aesthetic (schuhmann2022laion) (AES.) as our evaluation metrics. The outcomes of this quantitative assessment are detailed in Tab. 1 and Tab. 5. These tables reveal that integrating Self-NPO with the base model and its preference-optimized variants consistently improves the aesthetic quality of the generated images. Beyond presenting average scores, as shown in Fig. 7, we also compute the percentage of samples generated from the same prompt that attain a higher preference score. The results produced with Self-NPO markedly surpass those generated without it. For text-to-video generation, we provide detailed experimental comparisons and demonstrations in the supplementary material, where our method achieves clear performance improvements on the majority of metrics.
Qualitative comparison. We present the generated results in different scenarios in Fig. 2, Fig. 3, and Fig. 8, demonstrating that self-NPO consistently improves both the generation visual quality and fidelity.
User study. We evaluate generation quality across four areas: color and lighting, high-frequency details, low-frequency composition, and text-image alignment. In Color and Lighting, users assess whether the images feature natural, visually pleasing color schemes and lighting. For high-frequency details, users look at the sharpness and texture details, especially edges and fine elements. In low-frequency composition, they focus on the overall structure and balance of the image. Finally, for text-image alignment, users judge how well the generated image matches the input text. The user study uses prompts from Pickapic validation_unique dataset, with different models generating images based on the same random seed. Users choose between “No Preference” (Draws), “Self-NPO is better” (Wins), or “Self-NPO is worse” (Losses) for each pair of images. We collect responses from 9 volunteers, each evaluating 100 pairs of images generated by Dreamshaper, totaling 900 votes. Results, shown in Fig. 6, reveal that NPO significantly improves high-frequency details, enhances color and lighting preferences, and helps with composition. It also shows an improvement in text-image alignment.
Applications
@@
Controllable generation. Moreover, our approach seamlessly integrates with controllable generation techniques. By incorporating controllable plugins like T2I-Adapter (mou2024t2i), we significantly boost the model’s generation quality under layout-controlled conditions.
@@
Table 1: Quantitative performance comparison with stable diffusion v1-5 based models. means the metrics are copied from Diffusion-NPO. Other metrics are tested with official weights.
Method P.S. HPS. I.R. AES.
@@
SePPO + Self-NPO 21.73 30.28 0.6744 6.014